OpenStudy (anonymous):

PLEASE HELP WILL MEDAL!! given 2 points on the line, input the equation of the line in standard form Ax+By=C. Reduce all fractions to lowest terms (6,1) (5,4)

OpenStudy (xhawaiiangirlx):

Find the slope first Slope (M) - y2-y1 / x2-x1 = 5-0 / 1-8 = 5/-7 y= mx +b 5= 1(5/-7) +b ( y intercept) 5 - (5/-7) = b 5 5/7 = b Y= 5/-7x+5 5/7 standard form -5/-7x+y = 5 5/7 5-5/-7 = 4 7/7 - 5/-7 = 5 /7 The format is correct.

OpenStudy (xhawaiiangirlx):

that is an example

OpenStudy (xhawaiiangirlx):

use the same format to complete ur problem

OpenStudy (anonymous):

i dont know where to plug the points in though? im so confused idk why i just am .

OpenStudy (anonymous):

@XhawaiiangirlX